### v2

# Magnetic field spreading from stellar and galactic dynamos into the exterior

Abstract
The exteriors of stellar and galactic dynamos are usually modeled as a current-free potential field. A more realistic description might be that of a force-free magnetic field. Here, we suggest that, in the absence of outflows, neither of those reflect the actual behavior when the magnetic field spreads diffusively into a more poorly conducting turbulent exterior outside dynamo. In particular, we explain why the usual ordering of the dipole magnetic field being the most slowly decaying one is altered, and that the quadrupole can develop a toroidal component that decays even more slowly with radial distance. This behavior is best seen for spherical dynamo volumes and becomes more complicated for oblate ones. In either case, however, those fields are confined within a magnetosphere beyond which the field drops exponentially. The magnetosphere radius expands ballistically (i.e., linearly in time $t$) during the exponential growth phase of the dynamo, but diffusively proportional to $t^{1/2}$ during the saturated phase. We demonstrate that the Faraday displacement current, which plays a role in a vacuum, can safely be neglected in all cases. The superposition of magnetic fields from galaxies in the outskirts of the voids between galaxy clusters can therefore not explain the void magnetization of the intergalactic medium, reinforcing the conventional expectation that those fields are of primordial origin. For quadrupolar configurations, the synchrotron emission from the magnetosphere is found to be constant along concentric rings. Dipolar or the quadrupolar configurations display large-scale radial trends that are potentially distinguishable with existing radio telescopes.
Magnetic fields (994); Hydrodynamics (1963)
I Introduction
The magnetic fields in stars and galaxies can be explained by dynamo action converting the kinetic energy in the turbulence and the differential rotation into magnetic energy. Such systems are traditionally computed by assuming that the magnetic field outside the dynamo domain continues as a current-free potential field. Mathematically, this can be formulated as a suitable boundary condition applied on the outer radius of a spherical (Steenbeck & Krause, 1969) or an ellipsoidal domain (Stix, 1975). In cylindrical and several other coordinate systems, on the other hand, there exists no method to apply a potential field boundary condition. One possibility is then to adopt a perfectly conducting boundary condition (Meinel et al., 1990). However, a more physical procedure might be to embed the galaxy in a poorly conducting exterior (Elstner et al., 1990; Brandenburg et al., 1990), because this would approximate a vacuum in the limit of zero conductivity.
The assumption of a vacuum outside the dynamo was never thought to be more than just a mathematically convenient construct. However, in recent work of Garg et al. (2025), a vacuum was assumed to exist even in the far-field of galaxies. Their study, which does not involve outflows or winds of any type, spawned subsequent investigations that addressed the question what the nature of the magnetic field in the far-field of galaxies actually is (Seller & Sigl, 2025; Ghosh et al., 2026). The present paper is a follow-up of our earlier work (Ghosh et al., 2026). Here, we focus on a more detailed numerical study, approach in some cases the limit of extremely poor conductivity, and make contact with possible measurements of synchrotron emission.
It is commonly assumed that a more realistic boundary condition of a dynamo corresponds to an electrically conducting exterior, although not necessarily to a perfectly conducting one. In addition, if the density of the exterior is low and the exterior is at rest, the magnetic field would be force-free, i.e., the cross product of current density and magnetic field vanishes. Also in the context of plasma confinement configurations, force-free magnetic fields tend to be more realistic than the assumption of a vacuum (Freidberg, 2014). Since in MHD, the current density is proportional to the curl of the magnetic field, both fields must then be parallel to each other. Such fields would therefore be eigenfunctions of the curl operator and are also known as Chandrasekhar-Kendall functions (after the early paper by Chandrasekhar & Kendall, 1957).
Adopting a perfectly force-free magnetic field is generally problematic because it might topologically not be realizable and it would also take some time to establish such a field in the vast exterior of a dynamo. Therefore, to alleviate this problem, a commonly used approach in solar physics is the magneto-frictional one (Yang et al., 1986). Here, the Lorentz force is allowed to drive flows through a friction term. These flows tend to relax the field until the force is minimized. In a similar approach, we study here the spreading of galactic magnetic fields into intergalactic space by solving the mean-field momentum equation and modeling the dynamo exterior as a medium with a certain viscosity and electric conductivity. A vacuum would correspond to the limit of very small conductivity or very large magnetic diffusivity. A large effective magnetic diffusivity emerges in a turbulent environment when modelling the large-scale field as an average or mean field, $\overline{\bm{B}}$}, where the overbar denotes some suitable averaging. However, as we explain below, the diffusivity needed to reproduce the electric properties of a vacuum would be at least eight orders of magnitude larger than what can be explained by turbulence.
It is useful to represent the resulting magnetic field as a multipole expansion with terms proportional to $Y_{\ell}^{m}(\theta,\phi)$, where $\ell$ and $m$ are the spherical harmonic degree and order, respectively. For a vacuum, the magnetic field decays with radius $r$ like $r^{-(\ell+2)}$. The lowest multiple is the dipole with $\ell=1$, so the field decays like $r^{-3}$, while for a quadrupole, $\ell=2$, so the field decays like $r^{-4}$. In an electrically conducting environment, magnetic fields are expected to be close to force-free and can then decay differently. If the gas pressure is negligible, force-free magnetic fields obey ${\overline{\bm{J}}}\times{\overline{\bm{B}}}=0$, i.e., the current density ${\overline{\bm{J}}}={\bm{\nabla}}\times{\overline{\bm{B}}}/\mu_{0}$ with $\mu_{0}$ being the vacuum permeability, is parallel to $\overline{\bm{B}}$}. This means that ${\bm{\nabla}}\times{\overline{\bm{B}}}=\alpha_{\mathrm{ff}}{\overline{\bm{B}}}$, where $\alpha_{\mathrm{ff}}$ is a coefficient that must be constant along the magnetic field line, but it can vary perpendicular to it (Priest, 1982).
To explain the toroidal magnetic fields observed around solar-type stars, Bonanno (2016) and Bonanno & Del Sordo (2017) considered force-free magnetic fields in mean-field dynamo exteriors. They assumed the coronal magnetic field of those stars to be harmonic, but they did not consider the question how such magnetic field configurations can be achieved in a finite time. Here, we show that such fields can only be produced within a slowly expanding, but finite magnetosphere around the dynamo. This also has applications to galactic dynamos and the question whether such magnetic fields can explain the magnetic field in the voids between galaxy clusters (Garg et al., 2025; Seller & Sigl, 2025; Ghosh et al., 2026).
In Section II, we present the details of the model, allowing for different profiles of the effective magnetic diffusivity, as well as feedback from mean flows and/or a finite speed of light in some cases. Our results about the establishment of a time-dependent magnetosphere are discussed Section III. Using a similar model setup, Ghosh et al. (2026) showed that the growth of the magnetosphere changes from ballistic to diffusive once the dynamo saturates. They also demonstrated that a quadrupolar magnetic field in the magnetosphere decays with radius $r$ only like $r^{-2}$, which is slower than the $r^{-3}$ decay of a dipolar field—both in a vacuum and in a conducting exterior. Here we present a detailed analysis of these empirical findings. In Section IV, we discuss new observational constraints in terms of synchrotron radio emission, which go beyond the rotation measure diagnostics presented in Ghosh et al. (2026). Finally, we conclude in Section V.
II Our model
II.1 Basic equations
We solve the mean-field dynamo equations using spherical coordinates, $(r,\theta,\phi)$. We refer to earlier work by Jabbari et al. (2013) and Jakab & Brandenburg (2021) for a similar treatment. To quantify the difference between a poorly conducting exterior and a vacuum, we also include in some cases the Faraday displacement current. The detailed motivation for such a study along with first results was already presented by Ghosh et al. (2026).
The displacement current can play a role when the ordinary current density $\overline{\bm{J}}$} is small. We assume that the mean current density obeys Ohm’s law, which, in our case, takes the form
${\overline{\bm{J}}}=\sigma(\overline{\mbox{$E$}}{}+\overline{\bm{U}}\times{\overline{\bm{B}}}+\overline{\bm{u}\times\bm{b}}),$ (1)
where $\sigma$ is the microphysical conductivity, $\overline{\mbox{$E$}}{}$ is the mean electric field, $\overline{\bm{U}}$ is the mean velocity, $\overline{\bm{B}}$} is the mean magnetic field, and $\overline{\bm{u}\times\bm{b}}$ is the mean electromotive force resulting from the small-scale velocity and magnetic fields, $\bm{u}$ and $\bm{b}$, respectively. In its simplest form, the mean electromotive force can be written as
$\overline{\bm{u}\times\bm{b}}=\alpha{\overline{\bm{B}}}-\eta_{\rm turb}\,\mu_{0}{\overline{\bm{J}}},$ (2)
where $\alpha$ is a coefficient that emerges in the calculation of $\overline{\bm{u}\times\bm{b}}$ for helical turbulence (Moffatt, 1978). It can lead to the growth and sustenance of a mean magnetic field and is generally referred to as the $\alpha$ effect. Furthermore $\eta_{\rm turb}$ is the turbulent magnetic diffusivity.
It is convenient to define the microphysical magnetic diffusivity $\eta=(\mu_{0}\sigma)^{-1}$ along with the total (effective) magnetic diffusivity $\eta_{\rm eff}=\eta+\eta_{\rm turb}$. Ohm’s law for the mean current density can then be written as
$\mu_{0}{\overline{\bm{J}}}=(\overline{\mbox{$E$}}{}+\overline{\bm{U}}\times{\overline{\bm{B}}}+\alpha{\overline{\bm{B}}})/\eta_{\rm eff}.$ (3)
The term $\mu_{0}{\overline{\bm{J}}}$ enters as the second one on the right-hand side of the averaged Ampere–Maxwell equation,
$\frac{1}{c^{2}}\frac{\partial\overline{\mbox{$E$}}{}}{\partial t}={\bm{\nabla}}\times{\overline{\bm{B}}}-\mu_{0}{\overline{\bm{J}}}.$ (4)
where $c$ is the speed of light.
The full set of Maxwell equations can then be written as
$\frac{\partial{\overline{\bm{A}}}}{\partial t}=-\overline{\mbox{$E$}}{},\qquad{\overline{\bm{B}}}={\bm{\nabla}}\times{\overline{\bm{A}}},$ (5)
along with
$\left(1+\frac{\eta_{\rm eff}}{c^{2}}\frac{\partial}{\partial t}\right)\overline{\mbox{$E$}}{}=\overline{\bm{U}}\times{\overline{\bm{B}}}+\alpha{\overline{\bm{B}}}-\eta_{\rm eff}{\bm{\nabla}}\times{\overline{\bm{B}}},$ (6)
where $\overline{\bm{A}}$} is the mean magnetic vector potential, so the mean magnetic field is given by ${\overline{\bm{B}}}={\bm{\nabla}}\times{\overline{\bm{A}}}$, which satisfies the constraint ${\bm{\nabla}}\cdot{\overline{\bm{B}}}=0$. The mean charge density is given by $\overline{\rho}_{\mathrm{e}}=\epsilon_{0}{\bm{\nabla}}\cdot\overline{\mbox{$E$}}{}$, where $\epsilon_{0}=1/\mu_{0}c^{2}$ is the vacuum permittivity. Here, we have adopted the Weyl gauge, i.e., the electrostatic potential vanishes.
The Faraday displacement current, $\epsilon_{0}\partial\overline{\mbox{$E$}}{}/\partial t$, can be neglected when $\eta_{\rm eff}\ll\tau c^{2}$, where $\tau$ is a characteristic time scale. In that limit, the terms on the rhs of Equation (6) can directly be inserted on the rhs of Equation (5) to yield the familiar induction equation. We refer to Appendix A for a discussion of the dispersion relation of a dynamo with $\alpha$ effect and a finite value of $c$ for the simple Cartesian case in a periodic domain with constant coefficients and no feedback from mean gas motions, i.e., $\overline{\bm{U}}=0$. There, and later in the present paper, we identify $\tau$ with the light travel time, i.e., with $(ck)^{-1}$ or $R/c$, where $k$ is the wavenumber of the lowest dynamo mode in a periodic domain, and $R$ is the radius of the dynamo in an open domain. Thus, we have $\eta_{\rm eff}\ll c/k$ and $\eta_{\rm eff}\ll cR$ in these two cases.
It is important to stress that the nature of the basic equations changes from parabolic to hyperbolic when the displacement current is included. This also implies that the computational time step constraint changes from being quadratic to linear in the mesh spacing. This can significantly alleviate the computational restrictions of explicit schemes and can be particularly important in spherical computational domains near the center, where the mesh spacing becomes small.
To account for the mean gas motions driven by the magnetic field, we also solve the momentum equation assuming the gas to be isothermal, so the sound speed $c_{\rm s}$ is constant and the mean pressure $\overline{p}(x,z,t)$ is given by $\overline{p}=\overline{\rho}c_{\rm s}^{2}$. In contrast to Jakab & Brandenburg (2021), we ignore the effects of a Parker wind and therefore set the gravitational potential to zero.
The full system of equations is then to be complemented by those for $\overline{\rho}$ and $\overline{\bm{U}}$, i.e.,
$\frac{{\rm D}{}\ln\overline{\rho}}{{\rm D}{}t}=-{\bm{\nabla}}\cdot\overline{\bm{U}},$ (7)
$\frac{{\rm D}{}\overline{\bm{U}}}{{\rm D}{}t}=-c_{\rm s}^{2}{\bm{\nabla}}\ln\overline{\rho}+\left[{\overline{\bm{J}}}\times{\overline{\bm{B}}}+{\bm{\nabla}}\cdot(2\nu_{\rm eff}\overline{\rho}\overline{\mbox{${\mathsf{S}}$}}{})\right]/\overline{\rho},$ (8)
where ${\rm D}{}/{\rm D}{}t=\partial/\partial t+\overline{\bm{U}}\cdot{\bm{\nabla}}$ is the advective derivative and $\overline{\mbox{${\mathsf{S}}$}}{}$ is the rate-of-strain tensor of the mean flow with the components $\overline{\mathsf{S}}_{ij}=(\overline{U}_{i;j}+\overline{U}_{j;i})/2-\delta_{ij}{\bm{\nabla}}\bm{\cdot}\overline{\bm{U}}/3$. Here, semicolons denote covariant derivatives.
There are three mean-field parameters: the $\alpha$ effect, the turbulent (effective) viscosity $\nu_{\rm eff}$, and the turbulent (effective) magnetic diffusivity $\eta_{\rm eff}$. In addition to the macrophysical backreaction from the mean Lorentz force ${\overline{\bm{J}}}\times{\overline{\bm{B}}}$, we also allow for microphysical feedback in the form of $\alpha$ quenching (Ivanova & Ruzmaikin, 1977) and assume $\alpha$ to be proportional to a quenching factor
$Q({\overline{\bm{B}}})=1/(1+{\overline{\bm{B}}}^{2}/B_{\rm eq}^{2}),$ (9)
where $B_{\rm eq}$ is the equipartition field strength above which $\alpha$ begins to be affected by the feedback from the Lorentz force of the small-scale magnetic field.
II.2 Details of the model
In the following, we adopt the computational domain $r_{\mathrm{in}}\leq r\leq r_{\mathrm{out}}$ and $0\leq\theta\leq\pi/2$, and assume axisymmetry, i.e., $\partial/\partial\phi=0$. In Equation (3), we must specify spatial profiles for $\alpha$ and $\eta_{\rm eff}$. To model a localized dynamo-active region, we choose $\alpha$ to be finite around the origin and zero outside the dynamo region.
In Equation (3), we must specify spatial profiles for $\alpha$ and $\eta_{\rm eff}$. We mostly study spherical dynamo regions with a given radius $R$, but in some cases, we also allow the region to be oblate. In those cases, the disk thickness in the $z$ direction, i.e., along the axis, is denoted by $h$. To model a smooth transition, we choose an error function profile. In addition, $\alpha$ should be antisymmetric about $z=0$. This is accomplished by multiplying with an additional $z/h$ factor. Thus, we write the $\alpha$ effect in the form
$\alpha=\alpha_{0}\frac{z}{h}\,\frac{Q({\overline{\bm{B}}})}{2}\left[1-{\rm erf}\frac{\sqrt{(\varpi/R)^{2}+(z/h)^{2}}-1}{w/R}\right],$ (10)
where $\alpha_{0}$ is a constant, $\varpi=r\sin\theta$ is the cylindrical radius, $z=r\cos\theta$ is the height above the midplane, and $w$ is the width of the transition from the dynamo-active region to the exterior. For $h=R$, the dynamo region is spherical.
For $\eta_{\rm eff}$, we allow the effective magnetic diffusivity to be a larger value $\eta_{\rm eff}^{\mathrm{ext}}$ outside the dynamo region and equal to $\eta_{\rm eff}^{\mathrm{int}}\leq\eta_{\rm eff}^{\mathrm{ext}}$ within the dynamo region. We construct the profile for $\eta_{\rm eff}$ similarly to that of $\alpha$, except that it is symmetric about $z=0$ and there is no quenching factor. Thus, we choose the profile for $\eta_{\rm eff}$ to be given by
$\eta_{\rm eff}=\eta_{\rm eff}^{\mathrm{int}}+\frac{\eta_{\rm eff}^{\mathrm{ext}}-\eta_{\rm eff}^{\mathrm{int}}}{2}\left[1+{\rm erf}\frac{\sqrt{(\varpi/R)^{2}+(z/h)^{2}}-1}{w/R}\right].$ (11)
We also study some models where $\eta_{\rm eff}$ is uniform throughout the domain, i.e., $\eta_{\rm eff}^{\mathrm{ext}}=\eta_{\rm eff}^{\mathrm{int}}$, as well as models where $\eta_{\rm eff}$ grows gradually in the exterior like a power law. In those cases, we take
$\eta_{\rm eff}=\eta_{\rm eff}^{\mathrm{int}}\left(1+r/r_{0}\right)^{m},$ (12)
where $m>0$ is the exponent and the parameter $r_{0}$ determines how rapidly $\eta_{\rm eff}$ grows in the exterior.
To avoid the coordinate singularity at $r=0$, we assume $r_{\mathrm{in}}=0.1\,R$ in most cases. For the outer radius, we assume $r_{\mathrm{out}}=1000\,R$. We restrict ourselves to solving the equations in the first quadrant of the meridional plane, i.e., $0\leq\theta\leq\pi/2$. We adopt the following boundary conditions on $\theta=\pi/2$ (Jabbari et al., 2015):
$\frac{\partial\overline{A}_{r}}{\partial\theta}=\overline{A}_{\theta}=\frac{\partial\overline{A}_{\phi}}{\partial\theta}=0$ (13)
for dipolar (odd parity) fields, and
$\overline{A}_{r}=\frac{\partial\overline{A}_{\theta}}{\partial\theta}=\overline{A}_{\phi}=0$ (14)
for quadrupolar (even parity) fields. In the radial direction, we adopt a perfect conductor boundary condition on both ends, i.e., $\overline{A}_{\theta}=\overline{A}_{\phi}=0$. Analogous boundary conditions in both directions also apply to the components of $\overline{\mbox{$E$}}{}$.
In all our models, we use as initial condition weak Gaussian distributed noise for $\overline{\bm{A}}$}. It acts as a seed magnetic field that is needed to allow the dynamo to amplify this field exponentially once $\alpha_{0}$ is large enough. Owing to the $Q({\overline{\bm{B}}})$ factor, the growth gets quenched and $|{\overline{\bm{B}}}|$ reaches values close to $B_{\rm eq}$. The initial values of $\overline{\mbox{$E$}}{}$ and $\overline{\bm{U}}$ are set to zero and $\overline{\rho}$ is set to a constant reference value $\rho_{0}$. We recall that gravity is not included in our models.
The strength of the $\alpha$ effect is quantified by the dynamo number $C_{\alpha}=\alpha_{0}R/\eta_{\rm eff}^{\mathrm{int}}$. The diffusivity contrast is given by $C_{\eta}=\eta_{\rm eff}^{\mathrm{ext}}/\eta_{\rm eff}^{\mathrm{int}}$. Time is often expressed in terms of the diffusion time $\tau_{\mathrm{diff}}=R^{2}/\eta_{\rm eff}^{\mathrm{ext}}$ based on the exterior diffusivity. During the kinematic growth phase, the growth rate $\gamma$ of the magnetic field is characterized by the nondimensional number $C_{\gamma}=\gamma R^{2}/\eta_{\rm eff}^{\mathrm{int}}$. Unlike $C_{\alpha}$ and $C_{\eta}$, which are input parameters, $C_{\gamma}$ is an output parameter.
For all of our spherical simulations, we use $1024\times 32$ meshpoints in $r_{\mathrm{in}}\leq r\leq 1000\,R$ and $0\leq\theta\leq\pi/2$. We use the Pencil Code (Pencil Code Collaboration et al., 2021), where the inclusion of the displacement current is facilitated by the module special/disp_current.f90. Its mean-field implementation through the module magnetic/meanfield.f90. has been validated by reproducing $\gamma(k)$ from the dispersion relation for constant coefficients in the Cartesian case; see Appendix A.
As an empirical measure of the front speed in simulations with the displacement current included, we determine the radius $r_{0.01}$ where the compensated value of the latitudinally averaged magnetic field has dropped to 0.01 times the value inside the magnetosphere. We expect that $r_{0.01}/ct<1$ in all cases, which is indeed borne out by the simulations.
As already discussed in Ghosh et al. (2026), the front speed is constant during the kinematic growth phase. We call this growth ballistic. In the saturated state, the front speed decreases when the radius grows diffusively. Thus, for the front radius $r_{\ast}(t)$, we have
$r_{\ast}^{2}(t)=\left\{\begin{array}[]{ll}q_{\mathrm{ballistic}}\gamma\eta_{\rm eff}^{\mathrm{ext}}(t-t_{\ast})^{2}&\mbox{(kinematic phase),}\\ q_{\mathrm{diff}}\eta_{\rm eff}^{\mathrm{ext}}(t-t_{\ast})&\mbox{(saturated phase),}\end{array}\right$. (15)
where $q_{\mathrm{ballistic}}$ and $q_{\mathrm{diff}}$ are empirical coefficients of the order of unity and $t_{\ast}$ denotes the time when the corresponding growth behavior is fully established.
II.3 Application to galactic scales
In this work, we present our results in dimensionless form by expressing length in units of $R$ and time in units of $\tau_{\mathrm{diff}}\equiv R^{2}/\eta_{\rm eff}^{\mathrm{ext}}$. In a specific example, Ghosh et al. (2026) estimated $\eta_{\rm eff}^{\mathrm{ext}}=3000\,{\rm kpc}\,{\rm km\,s}^{-1}$. They considered three particular models which are also presented here as Runs A–C. They presented their results in dimensionful fashion. The underlying dimensionless time at the end of their runs, $t_{\mathrm{end}}/\tau_{\mathrm{diff}}$, was about 1400. Identifying $t_{\mathrm{end}}$ with the Hubble time of $14\,{\rm Gyr}$ yields $\tau_{\mathrm{diff}}=0.01\,{\rm Gyr}$. The unit of length is therefore $R=(\eta_{\rm eff}^{\mathrm{ext}}\tau_{\mathrm{diff}})^{1/2}\approx 5.5\,{\rm kpc}$. For one Hubble time, Ghosh et al. (2026) quoted a typical diffusion length of $200\,{\rm kpc}$; see their Table II for a case they referred to as cosmic web dynamics.
Table 1: Summary of runs presented in the paper.
Run sym $\overline{U}$ $cR/\eta_{\rm eff}$ $r_{\mathrm{in}}/R$ $h/R$ $C_{\alpha}$ $C_{\eta}$ $C_{\gamma}$ $B_{0}/B_{\rm eq}$ $q_{\rm diff}$aaA hyphen indicates that Equation (17) does not provide an appropriate fit. $r_{0.01}/ct$bbThe ratio $r_{0.01}/ct$ can only be evaluated when $c$ is finite. $n$ A D $\neq 0$ $\infty$ 0.1 1 25 50 63 0.7 4.9 3 B Q $\neq 0$ $\infty$ 0.1 1 25 50 66 0.4 2.5 2 C Q $\neq 0$ $\infty$ 0.1 1 25 10 70 0.4 2.5 2 D D 0 $\infty$ 0.1 1 25 1 74 0.7 4.9 3 E D 0 1 0.1 1 25 1 7.2 0.7 0.8 3 F D 0 0.5 0.1 1 25 1 3.7 0.7 0.6 3 G Q 0 $\infty$ 0.1 1 25 1 77 0.4 2.5 2 G’ Q 0 $\infty$ 0.1 1 25 1-100 68 0.4 2.5 2.16 G” Q 0 $\infty$ 0.1 1 25 1-100 68 0.4 2.5 2.23 H Q 0 1 0.1 1 25 1 8.2 0.4 0.7 2 I Q 0 0.5 0.1 1 25 1 4.2 0.4 0.6 2 J Q 0 $\infty$ 0.2 1 50 1 386 0.6 2.5 2 K Q 0 $\infty$ 0.1 0.5 50 1 348 0.2 2.5 2 L Q 0 $\infty$ 0.1 0.2 50 1 78 $2\times 10^{-5}$ 2.5 2
We reiterate that our intention is not to provide a realistic model of a galactic dynamo. Instead, we are interested in understanding the far-field behavior of a generic dynamo. Our results may therefore in principle also be applicable to stellar dynamos. We stress that differential rotation is not included either. All these simplifications are done to keep the results simple enough to be able to identify potentially generic behaviors that are expected to occur in the far-field of astrophysical dynamos in general.
III Results
III.1 Exponential growth phase
Let us first look at the initial spreading of a dynamo-generated magnetic field into $r>R$. We adopt an $\alpha$ effect in $r_{\mathrm{in}}\leq r\leq R$ with $C_{\alpha}=25$, and a 10 or 50-fold increased value of $\eta_{\rm eff}$ in the exterior; see Table LABEL:TSummary, where we present a summary of the runs discussed in this paper.
At early times, the magnetic field expands into $r>R$ and is found to fall off with distance like $e^{-\kappa r}$, where
$\kappa=(\gamma/\eta_{\rm turb})^{1/2}$ (16)
is the radial decay rate, $\gamma$ is the growth rate, and $\eta_{\rm turb}$ the turbulent magnetic diffusivity. This scaling is analogous to that for the skin effect, except that here we use an imaginary frequency, which leads to a non-oscillatory decay with increasing distance from the dynamo regime.
Since the magnetic field still increases exponentially like $e^{\gamma t}$, we have $B\sim e^{-\kappa r+\gamma t}=e^{-\kappa(r-c_{\mathrm{front}}t)}$, where $c_{\mathrm{front}}=\gamma/\kappa=(\gamma\eta_{\rm turb})^{1/2}$ is the speed of the front, where we have used Equation (16). After some initial transient, the front radius begins to increase linearly with time; see Figure 1. This scaling is similar to that for the propagation of fronts of diseases (Murray et al., 1986). It also applies to the turbulent propagation of reactive fronts in combustion when the reaction speed is faster than diffusion (Brandenburg et al., 2011).
Figure 1: Colorscale representation of $\ln|\bm{B}|$ vs $r$ and $t$ for Run B. Yellow (blue) shades denote large (small) fields. The dynamo operates in $0\leq r\leq 1$, as can be seen by the elevated field strength close to $r=0$. The left (right) dashed line corresponds to $q_{\mathrm{ballistic}}=1$ with $t_{\ast}/\tau=15$ (25).
Figure 2: Radial dependence of $\langle\bm{B}^{2}\rangle^{1/2}$ at times $t/\tau_{\mathrm{diff}}=30$, 35, and 50 (dotted lines), as well as $t/\tau_{\mathrm{diff}}=100$, 300, and 1000 (solid lines), for (a) a dipolar field with $C_{\eta}=50$ (Run A), (b) a quadrupolar field with $C_{\eta}=50$ (Run B), and (c) a quadrupolar field with $C_{\eta}=10$ (Run C). The asymptotic fall-offs $\propto r^{-3}$ for the dipolar field and $\propto r^{-2}$ for the quadrupolar fields are marked with dashed-dotted lines. The red lines denote the fit given by Equation (17) with the $q_{\rm diff}$ values listed in Table LABEL:TSummary.
III.2 Unconventional quadrupolar decay
In Figure 2, we plot the latitudinally averaged root-mean-square magnetic field for dipolar and quadrupolar boundary conditions. For the quadrupolar case, we also compare with a case with a five times smaller magnetic diffusivity in the exterior. These three models were also discussed in Ghosh et al. (2026). It turns out that, during the early kinematic growth phase, the magnetic field grows and spreads similarly regardless of its hemispheric parity or the value of the exterior magnetic diffusivity. After the magnetic field has saturated in the dynamo-active part, it continues to spread like $r^{-3}$ in the dipolar case, but like $r^{-2}$ in the two quadrupolar cases. This can clearly be seen from Figure 2. Note also that, during the kinematic phase, the fields decay more rapidly with radius than during the saturated phase. Indeed, this is a transient behavior where the magnetic field changes from the exponential radial decay, $\propto e^{-\kappa r}$, to a power law decay, $\propto r^{-n}$, in a shell around the dynamo, which then becomes the magnetosphere.
The radial powerlaw decay extends from the end of the dynamo-active part at $r=R$ to a radius $r_{\ast}(t)$ that we referred to as the front radius in Equation (15). It grows diffusively with time after some reference time $t_{\ast}$; see Figure 2 of Ghosh et al. (2026) for a plot of the spreading in radius and time. They referred to the sphere with radius $r_{\ast}(t)$ as the magnetosphere. For the nondimensional coefficient $q_{\rm diff}$, the empirically determined value is close to two—analogously to the law of Brownian diffusion (Einstein, 1905). In Figure 2, we also compare the modulus of the latitudinally averaged magnetic field for the last time step, with a fit of the form
$\overline{B}(r,t)=B_{0}\,\left(\frac{r}{1\,{\rm kpc}}\right)^{-n}\exp\{-\textstyle{\frac{1}{2}}[r/r_{\ast}(t)]^{2}\},$ (17)
where $B_{0}$ characterizes the magnetic field strength inside the magnetosphere and $n$ is the radial decay exponent. We find that the value of $q_{\rm diff}$ is around 4.9 for the dipolar configurations and around 2.5 for the quadrupolar configurations; see Table LABEL:TSummary. This demonstrates that the radius of the magnetosphere scales with the exterior magnetic diffusivity, $\eta_{\rm turb}^{\rm ext}$, and is also independent of the diffusivity contrast, $C_{\eta}$. Even the case $C_{\eta}=1$ results in the scaling expected for the value $\eta_{\rm turb}^{\rm ext}$, which is then also equal to $\eta_{\rm turb}^{\rm ext}$ in this case.
Following Ghosh et al. (2026), we determine the instantaneous front radius $r_{\ast}(t)$ as a weighted integral,
$\left.r_{\ast}(t)=\int r^{n+1}\overline{B}(r,t)\,dr\right/\!\int r^{n}\overline{B}(r,t)\,dr.$ (18)
The diffusive $r\propto t^{1/2}$ scaling is confirmed for all runs where $\eta_{\rm eff}={\rm const}{}=\eta_{\rm turb}^{\rm ext}$ in the exterior. However, for $\eta_{\rm eff}(r)$ profiles such as Equation (12), where $\eta_{\rm eff}\propto r^{m}$ for large radii, the growth is steeper: $\propto t^{0.8}$ for $m=1$ and $\propto t^{1.3}$ for $m=1.5$. This is shown in Figure 3, where we have used $r_{0}=10\,R$ in Equation (12). The shallow inversely quadratic radial fall-off of the mean magnetic field for the quadrupolar field is still approximately obeyed. In Appendix B, we demonstrate that an accurate measurement yields slightly steeper radial decays with $n\approx 2.16$ and 2.23 for models with $m=1$ and 3/2, respectively.
Figure 3: Comparison of $r_{\ast}(t)$ for Runs B (blue), G (black), G’ (red), and G” (orange).
To study the radial power laws for the dipolar and quadrupolar configurations in more detail, we now express the latitudinal dependence in terms of Legendre polynomials. An axisymmetric magnetic field can be written as (Moffatt, 1978; Parker, 1979; Krause & Rädler, 1980)
${\overline{\bm{B}}}=\hat{\bm{\phi}}\overline{B}_{\phi}+{\bm{\nabla}}\times\hat{\bm{\phi}}\overline{A}_{\phi},$ (19)
where $\overline{A}_{\phi}(r,\theta,t)$ and $\overline{B}_{\phi}(r,\theta,t)$ are independent of each other outside the dynamo region and characterize the poloidal and toroidal magnetic field decay in the radial direction. They can be expanded as [^1] In general, axisymmetric and nonaxisymmetric magnetic fields can be written as ${\overline{\bm{B}}}={\bm{\nabla}}\times\bm{r}T+{\bm{\nabla}}\times{\bm{\nabla}}\times\bm{r}S,$ (20) where $T(r,\theta,\phi,t)$ and $S(r,\theta,\phi,t)$ are the superpotentials for the three-dimensional toroidal and poloidal fields, respectively. In axisymmetry, this implies that $B_{\phi}=-\partial T/\partial\theta$, and if $T(r,\theta)=P_{\ell}(\cos\theta)$, then $B_{\phi}=-P_{\ell}^{1}(\cos\theta)$. Likewise, using $A_{\phi}=-\partial S/\partial\theta$, and if $S(r,\theta)=P_{\ell}(\cos\theta)$, then $A_{\phi}=-P_{\ell}^{1}(\cos\theta)$.
$\overline{A}_{\phi}=\sum a_{\ell}(r)\,P_{\ell}^{1}(\cos\theta),\quad\overline{B}_{\phi}=\sum b_{\ell}(r)\,P_{\ell}^{1}(\cos\theta).$ (21)
Dipolar (quadrupolar) magnetic fields have odd (even) $\ell$ in $a_{\ell}(r)$ and even (odd) $\ell$ in $b_{\ell}(r)$.
Figure 4: Radial dependence of (a) $a_{11}(r)$ and $b_{21}(r)$ for the dipole (Run A) and (b) $a_{21}(r)$ and $b_{11}(r)$ for the quadrupole (Run B). The asymptotic slopes are $a_{11}/(B_{\rm eq}R)\approx 0.052\,(r/R)^{-2}$ and $b_{21}/B_{\rm eq}\approx 0.3\,(r/R)^{-3}$ for the dipole and $b_{11}/B_{\rm eq}\approx 0.132\,(r/R)^{-2}$ and $a_{21}/(B_{\rm eq}R)\approx 0.037\,(r/R)^{-3}$ for the quadrupole, and are marked with dashed-dotted lines. For a vacuum field, $b_{11}$ would be zero. The red and blue lines give the scalings of $\overline{A}_{\phi}$ and $\overline{B}_{\phi}$, respectively, which are opposite for the dipolar and quadrupolar cases.
We compute the coefficients $a_{\ell}(r)$ and $b_{\ell}(r)$ from our simulations as
$a_{\ell}(r)=N_{\ell}\int\overline{A}_{\phi}(r,\theta)\,P_{\ell}^{1}(\cos\theta)\,{\rm d}{}\cos\theta,$ (22)
$b_{\ell}(r)=N_{\ell}\int\overline{B}_{\phi}(r,\theta)\,P_{\ell}^{1}(\cos\theta)\,{\rm d}{}\cos\theta,$ (23)
where $N_{\ell}=(2\ell+1)\,(\ell-1)!/(\ell+1)!$ is a normalization factor. In particular, $N_{1}=3/2$ and $N_{2}=5/6$. In Figure 4, we plot the radial dependences of $a_{11}(r)$ and $b_{21}(r)$ for the dipole and of $a_{21}(r)$ and $b_{11}(r)$ for the quadrupole. We clearly see the asymptotic slopes $a_{11}\propto r^{-2}$ and $b_{21}\propto r^{-3}$ for the dipole and $b_{11}\propto r^{-2}$ and $a_{21}\propto r^{-3}$ for the quadrupole.
Note that for the dipolar and quadrupolar solutions, the roles of $A_{\phi}$ and $B_{\phi}$ are interchanged; see Figure 4. In particular, the $r^{-2}$ behavior for $\overline{A}_{\phi}$ of the dipole is now seen for the $\overline{B}_{\phi}$ field of the quadrupole, and the $r^{-3}$ behavior for $\overline{B}_{\phi}$ of the dipole is now seen for the $\overline{A}_{\phi}$ field of the quadrupole. Thus, although the poloidal component for a quadrupolar configuration still decays like $r^{-4}$, the toroidal component $\overline{B}_{\phi}$ for this quadrupolar configuration decays like $\overline{A}_{\phi}$ for the dipolar configuration, i.e., like $r^{-2}$.
In the following, we discuss the individual components of the magnetic field in more detail. The two components of the poloidal field are given by $B_{r}={\rm D}_{\theta}{}A_{\phi}$ and $B_{\theta}=-{\rm D}_{r}{}A_{\phi}$, where ${\rm D}_{\theta}{}=r^{-1}\sin^{-1}\!\theta\,\partial_{\theta}(\sin\theta\,\cdot)$, and ${\rm D}_{r}{}=r^{-1}\partial_{r}(r\,\cdot)$ are differential operators.
Figure 5: Radial magnetic field profiles compensated (a) by $r^{3}$ for the dipolar case (Runs D–F), and (b) by $r^{2}$ for the quadrupolar case (Run G–I). For the solid lines, the displacement current is neglected (Runs D and G), while for the dashed and dotted lines it is included with $cR/\eta_{\rm eff}=1$ (Runs E and H) and 0.5 (Runs F and I), respectively. The red, blue, and black lines correspond to the times $t/\tau_{\mathrm{diff}}=20$, 200, and 800.
III.3 Lowest-order dipole
Employing our lowest-order dipole approximation, we find the two fields to be proportional to
$\overline{A}_{\phi}=a_{\mathrm{D}}\,r^{-2}\,P_{1}^{1}(\cos\theta),\quad\overline{B}_{\phi}=b_{\mathrm{D}}\,r^{-3}\,P_{2}^{1}(\cos\theta).$ (24)
Therefore, since $P_{1}^{1}(\cos\theta)=-\sin\theta$ and $P_{1}(\cos\theta)=\cos\theta$, we have
$\overline{B}_{r}=-2a_{\mathrm{D}}\,r^{-3}\,\cos\theta\equiv-2a_{\mathrm{D}}\,r^{-3}\,P_{1}(\cos\theta),$ (25)
$\overline{B}_{\theta}=-a_{\mathrm{D}}\,r^{-3}\,\sin\theta\equiv a_{\mathrm{D}}\,r^{-3}\,P_{1}^{1}(\cos\theta).$ (26)
For our model, we find $a_{\mathrm{D}}\approx 0.20\,B_{\rm eq}R^{3}$ and $b_{\mathrm{D}}\approx 0.60\,B_{\rm eq}R^{3}$.
III.4 Lowest-order quadrupole
Our lowest-order quadrupole is proportional to
$\overline{A}_{\phi}=a_{\mathrm{Q}}\,r^{-3}\,P_{2}^{1}(\cos\theta),\quad\overline{B}_{\phi}=b_{\mathrm{Q}}\,r^{-2}\,P_{1}^{1}(\cos\theta).$ (27)
Since $P_{2}^{1}(\cos\theta)=-3\sin\theta\cos\theta$ and $P_{2}(\cos\theta)=(3\cos^{2}\theta-1)/2$, we have
$\overline{B}_{r}=3a_{\mathrm{Q}}\,r^{-4}\,(1-3\cos^{2}\theta)\equiv-6a_{\mathrm{Q}}\,r^{-4}\,P_{2}(\cos\theta),$ (28)
$\overline{B}_{\theta}=-6a_{\mathrm{Q}}\,r^{-4}\,\sin\theta\cos\theta\equiv 2a_{\mathrm{Q}}\,r^{-4}\,P_{2}^{1}(\cos\theta).$ (29)
For our model, we find $a_{\mathrm{Q}}\approx 0.09\,B_{\rm eq}R^{4}$ and $b_{\mathrm{Q}}\approx 0.49\,B_{\rm eq}R^{2}$.
III.5 Current density profiles
Next, we compute $\mu_{0}{\overline{\bm{J}}}={\bm{\nabla}}\times{\overline{\bm{B}}}$, which is given by
$\mu_{0}{\overline{\bm{J}}}=\hat{\bm{\phi}}\mu_{0}\overline{J}_{\phi}+{\bm{\nabla}}\times\hat{\bm{\phi}}\overline{B}_{\phi},$ (30)
where
$\mu_{0}\overline{J}_{\phi}=-({\rm D}{}_{r}^{2}+{\cal D}_{\theta}^{2})\overline{A}_{\phi},$ (31)
is proportional to the toroidal current density. Again, ${\rm D}_{r}^{2}{}\equiv{\rm D}_{r}{}{\rm D}_{r}{}$ and ${\cal D}_{\theta}^{2}{}\equiv{\rm d}_{\theta}{}{\rm D}_{\theta}{}$ with ${\rm d}_{\theta}{}=r^{-1}\partial_{\theta}$ are convenient shorthands.
For the lowest-order dipole with $A_{\phi}=a_{\mathrm{D}}\,r^{-2}\,P_{1}^{1}(\cos\theta)$, we have
${\rm D}{}_{r}^{2}(-r^{-2}\sin\theta)=-2\,r^{-4}\sin\theta$ (32)
and
${\cal D}_{\theta}^{2}(-r^{-2}\sin\theta)=2\,r^{-4}\sin\theta,$ (33)
and therefore $\overline{J}_{\phi}=0$. Thus, we have $\mu_{0}{\overline{\bm{J}}}={\bm{\nabla}}\times\hat{\bm{\phi}}\overline{B}_{\phi}$, and therefore the Lorentz force, ${\overline{\bm{J}}}\times{\overline{\bm{B}}}$, is
${\overline{\bm{J}}}\times{\overline{\bm{B}}}=({\bm{\nabla}}\times\hat{\bm{\phi}}\overline{B}_{\phi})\times(\hat{\bm{\phi}}\overline{B}_{\phi}+{\bm{\nabla}}\times\hat{\bm{\phi}}\overline{A}_{\phi})/\mu_{0},$ (34)
and has only a purely toroidal component, whose divergence vanishes. Therefore, such a field is nearly force-free, except for the $\phi$ component, which only drives a rotational motion.
For the quadrupole with $\overline{A}_{\phi}\propto r^{-3}\sin\theta\cos\theta$, we have
${\rm D}{}_{r}^{2}r^{-3}=6\,r^{-5}$ (35)
and
${\cal D}_{\theta}^{2}\sin\theta\cos\theta=-6\,r^{-2}\sin\theta\cos\theta.$ (36)
Therefore, again, $\overline{J}_{\phi}=0$.
For the quadrupole, we have
$D^{2}\overline{B}_{\phi}=0,$ (37)
similarly to $\overline{A}_{\phi}$ for the dipolar field. However, unlike for $\overline{A}_{\phi}$, where $D^{2}\overline{B}_{\phi}=\mu_{0}\overline{J}_{\phi}$ is the toroidal current density, in the case of a quadrupolar field, $D^{2}\overline{B}_{\phi}$ has no physically obvious meaning. What is more relevant in a conductor is the Lorentz force, ${\overline{\bm{J}}}\times{\overline{\bm{B}}}$, which has poloidal and toroidal components. Its toroidal component coming from the poloidal components ${\overline{\bm{J}}}_{\mathrm{pol}}$ and ${\overline{\bm{B}}}_{\mathrm{pol}}$ is nonvanishing, but it only leads to driving an azimuthally differential rotation, and becomes extremely weak owing to its fast radial decay $\propto r^{-5}$. The poloidal Lorentz force has the components $^{\overline{\bm{J}}}_{\mathrm{pol}}\times\overline{B}_{\phi}\hat{\bm{\phi}}$ and $\hat{\bm{\phi}}D^{2}\overline{A}_{\phi}\times{\overline{\bm{B}}}_{\mathrm{pol}}/\mu_{0}$. Using $\mu_{0}{\overline{\bm{J}}}_{\mathrm{pol}}={\bm{\nabla}}\times\overline{B}_{\phi}\hat{\bm{\phi}}$ and $^{\overline{\bm{B}}}_{\mathrm{pol}}={\bm{\nabla}}\times\overline{A}_{\phi}\hat{\bm{\phi}}$, we have $({\bm{\nabla}}\times\overline{B}_{\phi}\hat{\bm{\phi}})\times\overline{B}_{\phi}\hat{\bm{\phi}}=-\varpi^{-1}{\bm{\nabla}}(\varpi\overline{B}_{\phi}^{2})$, and it leads to a radial force that falls off very rapidly like $r^{-5}$. Thus, except for a term $\overline{B}_{\phi}^{2}{\bm{\nabla}}\ln\varpi^{2}$, the Lorentz force can in principle be balanced by a pressure gradient. The residual $\overline{B}_{\phi}^{2}{\bm{\nabla}}\ln\varpi^{2}$ term, on the other hand, can be balanced by the centrifugal force. In that sense, our quadrupolar configuration is nearly force-free.
As already stated above, outside the dynamo region, $\overline{A}_{\phi}$ and $\overline{B}_{\phi}$ act independently of each other, and each component is subject to diffusion. Thus, unlike the dynamo interior, where $\overline{B}_{\phi}$ is coupled to $\overline{A}_{\phi}$ by induction effects such as differential rotation, it is here expanding just because of (turbulent) magnetic diffusion and can therefore adopt the same behavior as $\overline{A}_{\phi}$. In the present case, this means that for the quadrupolar configuration, $\overline{B}_{\phi}\propto r^{-2}$, just like the $\overline{A}_{\phi}\propto r^{-2}$ behavior for the dipolar configuration.
III.6 Effects of the displacement current
As we have seen from the Cartesian simulations in Appendix A, the inclusion of the displacement current can lead to a smaller growth rate $\gamma$; see Appendix A. It also can only decrease the front speed, $c_{\mathrm{front}}=\gamma/\kappa$. As a consequence, it turns out that the front speed is always less than the speed of light.
Figure 6: Radial magnetic field profiles for Runs J, K, and L. For Run J, we have $r_{\mathrm{in}}/R=0.2$ instead of 0.1, while for Runs K and L, we have $h/R=0.5$ and 0.2, respectively. The times are (a) $t/\tau_{\mathrm{diff}}=0.1$, 0.3, and 1.5 for the dotted lines and 10, 100, and 942 for the solid lines, (b) 0.1, 0.2, and 0.8 for the dotted lines and 10, 50, and 530 for the solid lines, (c) 0.2, 0.4, and 0.6 for the dotted lines and 500, 1200, 2700, and 5990 for the solid lines.
As alluded to in the introduction, the values of the magnetic diffusivity would need to be extremely large if one wanted to emulate the conditions of a vacuum. In vacuum, an electromagnetic wave propagates at the speed $c=300,000\,{\rm km\,s}^{-1}$, which implies a propagation distance $L=ct_{\mathrm{H}}\approx 5\,{\rm Gpc}$ in one Hubble time, $t_{\mathrm{H}}\approx 14\,{\rm Gyr}$. This corresponds to a reference diffusivity of about $\eta_{\mathrm{ref}}\equiv L^{2}/t_{\mathrm{H}}=c^{2}t_{\mathrm{H}}\approx 2\times 10^{12}\,{\rm kpc}\,{\rm km\,s}^{-1}$, which is nearly $10^{9}$ times larger than the value $\eta_{\rm turb}=3000\,{\rm kpc}\,{\rm km\,s}^{-1}$ used by Ghosh et al. (2026). Such large values are unrealistic. [^2] Even if the turbulent velocity was equal to the speed of light, it would require a correlation length of more than $70\,{\rm Gpc}$ to reach a turbulent diffusivity of $10^{12}\,{\rm kpc}\,{\rm km\,s}^{-1}$. Nevertheless, it is of interest to see how the spherical expansion of the magnetic field is affected by the finite speed of light in a medium of such an extremely large magnetic diffusivity. In the following, we consider the values $cR/\eta_{\mathrm{ref}}=1$ and 0.5. Again, these ratios are chosen to be unrealistically small and is done just to see the qualitative and quantitative effects of including the Faraday displacement current. We choose $C_{\eta}=1$, which, as demonstrated above, does not affect the radial expansion behavior compared to cases with larger values of $C_{\eta}$. The results for our models in spherical geometry are shown in Figure 5. For these models, we have neglected the evolution of the mean flow, i.e., we only solve Equations (5) and (6). It turns out that only at early times, the solution is noticeably affected by the finite speed of light—even for $cR/\eta_{\rm eff}=0.5$. In these cases, as explained in Section II.2, we characterize the front radius by the radius $r_{0.01}$, where the compensated magnetic field has dropped to 0.01 times the value inside the magnetosphere. We see from Table LABEL:TSummary that $r_{0.01}/ct<1$ is obeyed in all cases. This ratio varies between 0.8 and 0.6 for our dipolar configurations and between 0.7 and 0.6 for our quadrupolar configurations.
III.7 Effects of changing $r_{\mathrm{in}}$ and $h$
In Figure 6, we examine the effects of varying the ratios $r_{\mathrm{in}}/R$ and $h/R$ (Runs J–L). To ensure that the dynamo remains excited in spite of its smaller volume, we have here also increased the value of $C_{\alpha}$ from 25 to 50. We see that the radial fall-off is basically unchanged when increasing the value of $r_{\mathrm{in}}/R$ from 0.1 (Run G) to 0.2 (Run J). By contrast, when decreasing the value of $h/R$ from 1 (Run G) to $0.5$ (Run K) and $0.2$ (Run L), we see that the radial profiles become more complicated and develop radial oscillations. In addition, there is even an indication of a drop of the magnetic field in the wake of the front. To see this more clearly, we run the case with $h/R=0.2$ for a much longer time. The results suggests that the magnetic field takes the form of an azimuthal ring or at least a shell instead of a filled sphere. This complication would have obscured the otherwise simple radial powerlaw behaviors discussed above. This was also the reason why we have focussed our study mostly on the case $H/R=1$. It is important to note, however, that the position of the exponential drop at the edge of the magnetosphere is unchanged, as is seen by the unchanged value of the parameter $q_{\rm diff}$ listed in Table LABEL:TSummary and the fits presented in Figure 6 for the last time step.
IV Observational implications
To make contact with observations, Ghosh et al. (2026) computed the rotation measure (RM). Assuming a radial powerlaw dependence for the thermal electron density, $n_{\mathrm{th}}\propto r^{-s_{\mathrm{th}}}$, they found that RM falls off with radius like $r^{-2-s_{\mathrm{th}}}$ for the dipolar solution and like $r^{-1-s_{\mathrm{th}}}$ for the quadrupolar solution. Measuring such radial profiles can help to verify our expected scalings.
Another observational diagnostic is the resulting synchrotron emission. The synchrotron intensity $I(x,z,\lambda)$ and the complex polarization, $\mathcal{P}(x,z,\lambda)\equiv Q+{\rm i}U$ are obtained by line-of-sight integration. Here, $Q$ and $U$ are the Stokes parameters characterizing linear polarization and $\lambda$ is the wavelength.
Figure 7: Polarization vectors $(Q,U)$ superimposed on a color scale representation of the logarithmic intensity $I(x,z)$ (in Jy/sr) for Runs A (dipolar configuration, left) and B (quadrupolar configuration, right).
To get an idea about the basic properties of $I$, $Q$, and $U$, we adapt the expression for the radial scaling of the magnetic field given by Equation (17) to the lowest order dipole and quadrupole configurations of Sections III.3 and III.4. We then compute the components
$\overline{B}_{x}=\sin\theta\cos\phi\,\overline{B}_{r}+\cos\theta\cos\phi\,\overline{B}_{\theta}-\sin\phi\,\overline{B}_{\phi},$ (38)
$\overline{B}_{z}=\cos\theta\,\overline{B}_{r}-\sin\theta\,\overline{B}_{\theta},$ (39)
in Cartesian coordinates where $\phi$ is the azimuthal angle and $\theta$ is the colatitude. The synchrotron intensity $I(x,z)$ projected onto the $xz$ plane is given by
$I(x,z,\lambda)=\int_{-L}^{L}\epsilon(x,y^{\prime},z,\lambda)\,{\rm d}{}y^{\prime},$ (40)
where $\epsilon\propto n_{\mathrm{CR}}\overline{B}_{\perp}^{(\varsigma+1)/2}\lambda^{(\varsigma-1)/2}$ is the emissivity, $n_{\mathrm{CR}}$ is the cosmic ray density, $\lambda$ is the wavelength, and $\varsigma$ is the spectral index of the momentum distribution (Ginzburg & Syrovatskii, 1965). Realistic values for $\varsigma$ are in the range 2.8–3.2 as a result of the combination of the continuous injection of secondary electrons (by hadronic collisions), the injection of primary electrons by supernova-driven shocks, and the modulation by energy-dependent diffusion (e.g., Beck, 1991; Werhahn et al., 2021; Armillotta et al., 2025). Here, for simplicity, we use $\varsigma=3$, because then we have $\epsilon\propto\overline{B}_{\perp}^{2}$ and the complex intrinsic polarization is simply proportional to $\mathcal{B}_{\perp}^{2}$, where $\mathcal{B}_{\perp}=\overline{B}_{x}+{\rm i}\overline{B}_{z}$ is the complex magnetic field in the plane of the sky; see Brandenburg & Stepanov (2014) for earlier work utilizing this formalism.
We ignore here the effects of Faraday rotation, which would have introduced a factor $\exp(2{\rm i}\phi\lambda^{2})$ under the integral of $\mathcal{P}$, on the basis that we showed already in Ghosh et al. (2026) that this is negligible but in the galaxy core. Here, $\phi(x,y,z)=-K\int_{0}^{y}(n_{\mathrm{th}}\overline{B}_{y})(x,y^{\prime},z)\,{\rm d}{}y^{\prime}$ is the Faraday depth and $K$ is a known constant. Neglecting Faraday rotation corresponds to taking the limit $\lambda\to 0$.
In Figure 7, we present synthetic maps of the synchrotron intensity $I(x,z)$ for the same dipolar and quadrupolar configurations as in Ghosh et al. (2026), corresponding to Runs A and B, respectively. For $n_{\mathrm{CR}}$, we assume for simplicity a steady radial profile similarly to $n_{\mathrm{th}}$, which we write here as $n_{\mathrm{CR}}\propto r^{-s_{\mathrm{CR}}}$, analogously to the expression for $n_{\mathrm{th}}$ used by Ghosh et al. (2026). Here, we simply assumed $s_{\mathrm{CR}}=1$, although it should be noticed that radiative losses may complicate this trend in reality (e.g., Peron et al., 2021; Pfrommer et al., 2022; Armillotta et al., 2025). To have a quantitative estimate of the magnetospheric emission, we used (Rybicki & Lightman, 1986, see also Appendix C)
$\epsilon\,{\rm d}{}y\approx\frac{5\,\mu\mathrm{Jy}}{\mathrm{sr}}\left(\frac{n_{\mathrm{CR}}^{(100\,{\rm kpc})}}{10^{-7}\,{\rm cm}^{-3}}\right)\left(\frac{B_{\rm eq}}{1\,\mu{\rm G}}\right)^{2}\left(\frac{\lambda}{2\,{\rm m}}\right)\left(\frac{{\rm d}{}y}{1\,{\rm kpc}}\right),$ (41)
where $n_{\mathrm{CR}}^{(100\,{\rm kpc})}$ is the cosmic ray density at a galactocentric distance of $100\,{\rm kpc}$, assuming that $s_{\mathrm{CR}}=1$ may still be valid at that distance, which depends on the long-term integrated dispersal of cosmic rays around galaxies and in large-scale structures (see e.g. Vazza et al., 2025, for recent numerical studies).
In Figure 7, we overlay the polarization vectors $(Q,U)$. For a powerlaw dependence of the cosmic ray number density, $n_{\mathrm{CR}}\propto r^{-s_{\mathrm{CR}}}$, assuming $\overline{B}_{\perp}\propto r^{-n}$ with $n=2$ for quadrupoles and $n=3$ for dipoles, the intrinsic emissivity becomes proportional to $n_{\mathrm{CR}}\overline{B}_{\perp}^{2}\propto r^{-s_{\mathrm{CR}}-2n}$; see Equation (17). Line-of-sight integration adds another power, yielding $r^{-s_{\mathrm{CR}}-2n+1}$. Thus, for quadrupolar (dipolar) configurations, we find $I\propto r^{-3-s_{\mathrm{CR}}}$ ($\propto r^{-5-s_{\mathrm{CR}}}$).
We see that the polarization vectors are predominantly in the $x$ direction. This is a commonly observed feature related to the dominance of the toroidal magnetic field; see, e.g., Brandenburg et al. (1993) and Elstner et al. (1995) for earlier work where, unlike the present case, a radial wind was included. It is only for the dipolar configuration that the toroidal field changes sign at the midplane. This causes clear departures from the otherwise nearly perfect alignment of the polarization vectors with the $x$ direction.
As expected, the radial fall-off of the synchrotron intensity is still rather steep—even for the quadrupolar configuration. These models of intensity distributions may have already been observed in real galaxies as they produce differences in principle detectable with current radio instruments. For example, an $8$ hour integration with the LOFAR High Band Antenna (HBA) at $\approx 150\,{\rm MHz}$ typically reaches a noise level of $3\sigma_{\rm rms}\approx 1\,{\mu\rm Jy}/\rm arcsec^{2}=2.35\cdot 10^{-17}\,{\rm Jy}/{\rm sr}$. Based on Figure 7, a fraction of the expected radio emission should be detectable in both cases by radio telescopes, yet in the dipolar configuration the detectable region should be confined within a $\leq 200\rm~kpc$ radius, while in the quadrupolar case a much larger region should be detectable. However, for a quantitative prediction one also need to fix the radial distribution of cosmic rays, which can only roughly be guessed at this stage, while in real galaxies it must affected by the presence of large-scale outflows, especially at larger radii where the inertia of cosmic rays is expected to become dynamically dominant (e.g., Chiu et al., 2024). Interestingly however, the differences in the orientation of polarization vectors are rather independent on the distribution of relativistic electrons, and they can provide a more immediate way of testing the large distribution of magnetic fields. Intriguingly, preliminary work using LOFAR observations may already have yielded results that are possibly consistent with our predictions (Oei, 2025).
V Conclusions
We have studied the radial spreading of a dynamo-generated magnetic field into its electrically conducting surroundings. Unlike conventional dynamo theory, where the exterior field is assumed to be a current-free potential field, ours has electric currents and is only approximately force-free. As already reported by Ghosh et al. (2026), the quadrupolar field has a toroidal component, that decays with radial distance only like $r^{-2}$, which is slower than the $r^{-3}$ decay of the poloidal component for the dipolar configuration in the far-field of the dynamo. We are not aware of earlier reports of such a slow radial magnetic field decay.
An important difference compared to earlier work on force-free magnetic fields in the dynamo exterior, lies in the fact that we study the time-dependent case with direct numerical simulations, where a nearly force-free field is only gradually building up in an electrically conducting environment. As far as the formulation of the effective boundary condition of the dynamo is concerned, which was the main purpose of considering force-free exteriors (Bonanno, 2016; Bonanno & Del Sordo, 2017), the limited time available in establishing the exterior field may not be very important on such short length scales. However, the slower radial decay and the sharp exponential fall-off marking the end of the magnetosphere may be of observational interest (Ghosh et al., 2026) and might be tested with the new generation of radio observatories (e.g., Ghasemi-Nodehi et al., 2022). This feature could be used to distinguish the field from a dipolar one, on the one hand, and from one governed by a wind, on the other. Such winds have indeed frequently been discussed in simulations (Aramburo-García et al., 2021, 2022; Bondarenko et al., 2022), including cases with a dynamo coupled either an imposed wind (Brandenburg et al., 1993) or to a dynamically sustained one (Perri et al., 2021; Jakab & Brandenburg, 2021).
In the work of Brandenburg et al. (1993), maps of synchrotron emission were presented, but the displayed results do not allow us to determine whether the radial fall-off is characterized by a power law. Jakab & Brandenburg (2021) did not compute synchrotron emission, but they presented their results for $\overline{B}_{\phi}$ by compensating them by $r^{2}$. This shows that, although their field configuration was dipolar, the presence of a wind alone can also lead to a slow radial decay of $\overline{B}_{\phi}$.
In the present cases, an important clue in understanding the origin of the $r^{-2}$ decay of $\overline{B}_{\phi}$ comes from the correspondence between the pair $(\overline{B}_{\phi},\overline{A}_{\phi})$ for a quadrupolar configuration and the pair $(\overline{A}_{\phi},\overline{B}_{\phi})$ for a dipolar configuration. In both cases, the residual decays are of the form
$\left.\begin{array}[]{l}(\overline{B}_{\phi},\overline{A}_{\phi})\cr(\overline{A}_{\phi},\overline{B}_{\phi})\end{array}\right\}=\left[r^{-2}\,\,P_{1}^{1}(\cos\theta),\;r^{-3}\,\,P_{2}^{1}(\cos\theta)\right],$ (42)
and their evolution is simply the result of diffusion. Unlike the action of $\alpha$ and $\Omega$ effects in the dynamo interior, diffusion does not couple $\overline{A}_{\phi}$ and $\overline{B}_{\phi}$. This is why the aforementioned correspondence is possible.
While these results are also borne out in cases where the displacement current is included, we have shown that it can safely be ignored in all cases of interest. Its main role lies in limiting the expansion speed in poorly conducting media rather than facilitating a radial expansion.
It is also interesting to note that the far-field magnetic field configurations are nearly force-free in the sense that the resulting Lorentz force can be balanced by pressure gradient and centrifugal forces. However, this aspect may well be subdominant because the radial fall-off of the Lorentz force is rapid and our results are essentially independent of whether the momentum equation is included in our simulations.
In the present work, we have made a number of simplifying assumptions that has helped producing clean results. In particular, the assumption of a spherical rather than an oblate dynamo geometry have helped in producing powerlaw behaviors of the radial magnetic field decay. It is possible that at large distances from the dynamo region, this assumption becomes less critical. However, the present results for oblate dynamo geometries rather suggest that magnetosphere may take the form of a ring or a spherical shell. The inclusion of differential rotation in the dynamo region would complicate the results further. Preliminary studies have shown that the $r^{-2}$ fall-off for the quadrupolar configuration might become steeper. It is unclear, however, how this changes at larger distances. Further modifications are expected if we allowed for a radial wind to develop. In that case, however, it is conceivable that one might find results that depend on the activity of the galaxy in driving such a wind. Nevertheless, a slow radial fall-off for quadrupolar fields may still survive under certain more realistic conditions. As explained above, the resulting synchrotron emission may be of observational interest. However, as already stressed by Ghosh et al. (2026), the radii of galactic magnetospheres are not affected by this and remain prohibitively small in view of proposals by Garg et al. (2025) to explain the lower limits of void magnetization by the superposition of dipoles from the far-field of galaxies.
We thank the referee for a constructive review. We acknowledge stimulating discussions with Ruth Durrer, Deepen Garg, and Jennifer Schober. This research was supported in part by the Swedish Research Council (Vetenskapsrådet) under grant No. 2025-05957, the National Science Foundation under grants No. NSF AST-2307698, AST-2408411, and NASA Award 80NSSC22K0825. OG acknowledges support from a Wallenberg Academy Fellowship (PI Azadeh Fattahi). Additionally, OG is supported by the Swedish Research Council (Vetenskapsrådet) under contracts 2022-04283 and 2019-02337 and additionally by the Göran Gustafsson Foundation for Research in Natural Sciences and Medicine. AN is partially supported by the French National Research Agency (ANR) grant ANR-24-CE31-4686. FV has been partially supported by Fondazione Cariplo and Fondazione CDP, through grant n Rif: 2022-2088 CUP J33C22004310003 for the “BREAKTHRU” project. We acknowledge the allocation of computing resources provided by the Swedish National Allocations Committee at the Center for Parallel Computers at the Royal Institute of Technology in Stockholm. Software and Data Availability. The source code used for the simulations of this study, the Pencil Code (Pencil Code Collaboration et al., 2021), is freely available on https://github.com/pencil-code. The simulation setups and corresponding input and reduced output data are freely available on http://doi.org/10.5281/zenodo.18564321 (catalog http://doi.org/10.5281/zenodo.18564321).
Appendix A Dispersion relation for a dynamo with constant coefficients
To illuminate the mean-field dynamo behavior for a finite speed of light and to provide a benchmark for testing purposes, we discuss here the case of an $\alpha^{2}$ dynamo with constant coefficients. Seeking solutions proportional to $\exp(\gamma t+{\rm i}\bm{k}\cdot\bm{x})$, the dispersion relation $\gamma(\bm{k})$ can be obtained by solving for the roots of
$\left[\gamma+\eta_{\rm eff}k^{2}\left(1+\frac{\gamma^{2}}{c^{2}k^{2}}\right)\right]^{2}-\alpha^{2}k^{2}=0.$ (A1)
In the limit $c\to\infty$, the dispersion relation agrees with the conventional one,
$\gamma=\pm|\alpha|k-\eta_{\rm eff}k^{2}.$ (A2)
Figure 8 shows $\gamma(\alpha)$ for a fixed value of $k$ and three values of $c/\eta_{\rm eff}k$. The dependence matches the conventional one for $c\to\infty$. We see that the effect of a finite speed of light is to lower the value of $\gamma$.
Figure 8: Growth rate versus dynamo number $\alpha/\eta_{\rm eff}k$ for a fixed value of $k$ and $c/\eta_{\rm eff}k=0.5$ (red), 1 (green), and 2 (blue). The gray areas mark the regions of no growth. The conventional dependence for $c\to\infty$ corresponds to the black line. The black plus signs refer to data points obtained with the Pencil Code.
For an exponentially growing magnetic field, the spatial spreading can be described by writing ${\overline{\bm{B}}}\propto e^{-\kappa(z-c_{\mathrm{front}}t)}$, where $c_{\mathrm{front}}$ is the front speed, $\kappa c_{\mathrm{front}}=\gamma$ is the growth rate, and $\kappa=k$. This relation is only valid outside the dynamo-active region, where $\alpha=0$. Since the effect of a finite speed of light is to lower the value of $\gamma$, this also lowers the front speed $\gamma/k$. For the conventional dispersion relation, the front speed is
$c_{\mathrm{front}}=\gamma/k=|\alpha|-\eta_{\rm eff}k.$ (A3)
For finite values of $c$, the front speed is lowered, except for the marginal point when $k$ exceeds the critical value below which dynamo action is possible. This is shown in Figure 9 for three finite values of $c/\eta_{\rm eff}k$. We see that the line for $c/\eta_{\rm eff}k=2$ is already rather close to the case $c\to\infty$.
Figure 9: Front speed $\gamma/k$ (in units of $c$) versus $k$ (in units of $c/\eta_{\rm eff}$) for $\alpha/c=1$ and $c/\eta_{\rm eff}k=0.5$ (red), 1 (green), and 2 (blue). As in Figure 8, the gray areas mark the regions of no growth. The conventional line for $c\to\infty$ corresponds to the black line.
As expected, the front speed never exceeds the speed of light. By contrast, however, the input parameters $\alpha$ and $\eta_{\rm eff}k$, which also have the dimension of a speed, can exceed the value of $c$. Whether or not this can physically be realized is questionable.
Appendix B Power law growth of $\eta_{\rm eff}$
For the quadrupolar case, the inversely quadratic radial fall-off is still approximately obeyed even when $\eta_{\rm eff}$ is not constant for $r>R$. This is demonstrated Figure 10, where we compare the mean radial magnetic field profiles for Runs G, G’, and G”. The slopes do not depart markedly from the $r^{-2}$ decay. For a more quantitative demonstration, we show for Runs G’ and G” also the profiles compensated by $r^{+m}$ with $m=2.16$ and with $m=2.23$, respectively. This confirms that the decay exponents given in Table LABEL:TSummary are accurately determined.
Figure 10: (a) Radial magnetic field profiles for the quadrupolar Runs G, G’, and G”. Runs G’ and G” have a radially increasing $\eta_{\rm eff}(r)$ profile. In panels (b) and (c), the mean magnetic field profiles are compensated by $r^{2.16}$ and $r^{2.23}$, respectively. The red, blue, and black lines refer to the times $t/\tau_{\mathrm{diff}}=100$, 250, and 600 for Run G’, and $t/\tau_{\mathrm{diff}}=50$, 100, and 250 for Run G”.
Appendix C Derivation of Equation (41)
The purpose of this appendix is to provide the detailed derivation of Equation (41). Following Rybicki & Lightman (1986), for a powerlaw distribution of CRs such as $N(E)dE=\mathrm{C}E^{-\varsigma}dE,\quad E_{\rm min}<E<E_{\rm max}$ with a particle distribution index $\varsigma$ corresponding to the spectral index $s$ as $s=(\varsigma-1)/2$ such that $P_{\rm tot}(\omega)\propto\omega^{-(\varsigma-1)/2}$, normalization can be determined from the cosmic ray number density
$n_{\mathrm{CR}}=C\frac{E_{\max}^{1-\varsigma}-E_{\min}^{1-\varsigma}}{1-\varsigma},$ (C1)
for $E_{\rm min}<E<E_{\rm max}$ known as the band-limited case. However, if $E_{\rm max}\rightarrow\infty$, normalization $C$ depends only on the threshold energy $E_{\rm min}$ for $\varsigma>1$ as
$C=n_{\mathrm{CR}}(\varsigma-1)E_{\min}^{\varsigma-1}.$ (C2)
The total synchrotron power per unit volume per unit frequency irradiated by such a population of CR can be written as
$P_{\mathrm{tot}}(\nu)$ $=\frac{\sqrt{3}\,q^{3}\mathrm{C}\,B\sin\alpha}{mc^{2}(\varsigma+1)}\,C_{2}\,\left(\frac{2\pi mc\nu}{3qB\sin\alpha}\right)^{\frac{1-\varsigma}{2}},$ (C3)
Based on this, the optically thick surface brightness can be expressed in general as
$I_{\nu}\propto n_{\mathrm{C}R}\left(>E_{\min}\right)E_{\min}^{\varsigma-1}B^{(1+\varsigma)/2}L\nu^{(1-\varsigma)/2}.$ (C4)
Taking equipartition value of the magnetic field $B=B_{\rm eq}$, averaging over the pitch angle factor $\langle\sin^{(1+\varsigma)/2}\alpha\rangle$, and finally setting the index $\varsigma=3$, we obtain an optically thick surface brightness of the form
$I_{\nu}$ $\approx 5\times 10^{-6}\,\mathrm{Jy}/\mathrm{sr}\left(\frac{n_{\mathrm{CR}}}{10^{-7}\,\mathrm{cm}^{-3}}\right)\left(\frac{E_{\min}}{100\,\mathrm{MeV}}\right)^{2}$ (C5)
$\times\left(\frac{B}{1\,\mu\mathrm{G}}\right)^{2}\left(\frac{L}{1\,\mathrm{kpc}}\right)\left(\frac{\nu}{150\,\mathrm{MHz}}\right)^{-1}.$
Replacing $(\nu/150\,\mathrm{MHz})^{-1}$ by $(\lambda/2\,\mathrm{m})$ and using $E_{\min}=100\,\mathrm{MeV}$, we arrive at Equation (41).